# A strong–weak duality for the 1d long-range Ising model

### v1

@@ 1 Introduction and motivation
Thus, while the generalized free field $\chi$, with $\Delta_{\chi}=(1+s)/2$, coupled to the topological operator $\sigma$, with $\Delta_{\sigma}=0$, via $\sigma\chi$ correctly identifies the crossover location at $s^{\star}=1$, it fails to provide a correct description of the critical theory. Indeed, $\sigma$ never becomes a genuine local field, it does not acquire an anomalous dimension, and the relevant operator controlling the deviation from criticality is also absent. A different perspective is therefore needed in order to capture the full operator spectrum and scaling behavior near the crossover to short range in one dimension.
@@
  • For all $s\leq 1$, we propose a field theory that is weakly coupled near the crossover, exactly solvable at $s=1$, that reduces to the Kondo model at $s=1$, and that upon perturbative expansion reproduces the AYK model. We derive several predictions for CFT data at next-to-leading or next-to-next-to-leading order in perturbation theory, thus showing that such model provides not only a conceptual framework to recast the AYK model as a field theory, but also a computational tool to treat the near-crossover regime with systematic perturbative methods.
@@ 2.2.1 Unitary 1d CFTs with parity symmetry: the bootstrap definition
The lowest-weight state in $D_{\Delta_{i},J_{i}}$ corresponds to a local primary operator $\phi_{i}$ at $x=0$. $\Delta_{i}$ is the scaling dimension of $\phi_{i}$, and $J_{i}\in\{0,1\}$ its parity, so that $x\mapsto-x$ acts as $\phi_{i}(0)\mapsto(-1)^{J_{i}}\phi_{i}(0).$ The sequence of ordered pairs $((\Delta_{i},J_{i}))_{i=0}^{\infty}$ is called the spectrum of the theory. Reflection positivity implies $\Delta_{i}\geq 0$. We have $\Delta_{0}=0$, $J_{0}=0$, corresponding to the vacuum (in other words, $\phi_{0}$ is identity operator $\mathds{1}$). Correlation functions of $\phi_{i}$ are invariant under $\mathrm{PGL}_{2}(\mathbb{R})$, which maps point $x$ to $\frac{ax+b}{cx+d}$ and the field $\phi$ to $\phi^{\prime}$, given by [^5] Here |cx+a||-cx+a| is defined using a representative $g\in\mathrm{GL}_{2}(\mathbb{R})$ of $g]\in\mathrm{PGL}_{2}(\mathbb{R})$ with $|\mathrm{det}(g)|=1$.
@@
holding for all $i,j,k,\ell\in\mathbb{Z}_{>0}$ and all $0<z<1$. Here $G^{\Delta_{i},\Delta_{j},\Delta_{k},\Delta_{\ell}}_{\Delta_{m}}(z)$ are the 1d conformal blocks, first computed in [41], taking the form
@@ 2.3 Anderson-Yuval-Kosterlitz model
Figure 3: The RG flow for the beta functions in (2.21), at $s=1$ and $s=0.9$. Arrows point towards the IR. At $s=0$, the $g=0$ axis is a line of fixes points, but we have emphasized in blue and red the special points $\mathcal{J}=0$ and $\mathcal{J}=1$, respectively. The red lines separate the diagram into three or four regions: region II is the low-temperature broken phase ($g=0$ in the IR); region IIII is the high-temperature symmetric phase ($g\sim O(1)$ in the IR); regions $III$ and $IV$ are inaccessible to the LRI. The latter indeed corresponds to a one-dimensional subspace such as the dashed blue line. The intersection of such line with the red line defines the critical coupling $\mathcal{J}_{c}$ (i.e. critical temperature at fixed bond strength $J$, if $\mathcal{J}=J/T$).
@@ 2.4 Insights from the 1d short-range Ising model
Another important insight for developing a theory of the LRI model near $s=1$ comes from scrutinizing the possible physical content of the theory at $s>1$. As reviewed above, for $d>1$ and $s\geq s^{*}$, the LRI crosses to the SRI universality plus a decoupled GFF sector [10, 9]. What changes when $d=1$? As we argue below, the physics is bound to be not so different from that of the classical 1d SRI model, indeed:
@@
Given the facts we just recalled about the SRI, we would expect that at $s\geq 1$ the critical LRI, with classical Hamiltonian (1.1), would be described by the degrees of freedom of a $\mathbb{C}^{2}$ Hilbert space, with trivial correlators. However, as in $d>1$, this cannot be the end of the story, because it would imply that at the crossover some operators would disappear from the spectrum. For example, by continuity of spectrum, at $s=1$ we should find two $\mathbb{Z}_{2}$-odd operators with opposite parity and of dimension exactly equal to one, corresponding to $\varphi^{3}$ and $\partial\varphi$ in the $\varphi^{4}$ description (see Section 2.2), but clearly the SRI alone cannot provide such operators.101010 Following [9, 10], we could try to account for the missing operators by adding to the 1d SRI a GFF $\chi$ of scaling dimension $\Delta_{\chi}=(1+s)/2$, that could be identified with $\varphi^{3}$, and that should be decoupled from the spin degree of freedom at $s\geq 1$. However, we would still be missing the parity-odd operator corresponding to $\partial\varphi$, as it would be unclear how to introduce derivatives of $\hat{\sigma}_{3}$. Moreover, it would be unclear how to couple the two sectors at $s<1$ in order to flow to the LRI CFT: if for example, combining insights from the above reminder of 1d SRI and from [9, 10], we introduce the defect operator $\operatorname{tr}[\exp(h\,\int{\rm d}x\,\hat{\sigma}_{3}\,\chi(x))]$ in correlators over the $\mathbb{C}^{2}+$GFF theory, we obtain a trivial result, as the trace vanishes.
@@ 3.1 Compact generalized free fields with negative scaling dimension
As we review in Appendix A, for $1<\Delta_{\phi}\leq 0$, the noncompact 1d GFF can be rigorously defined [57] as a probability measure on the space of distributions $\mathbb{R}\rightarrow\mathbb{R}$ defined modulo an additive constant. In fact, the resulting random distribution is almost surely a continuous function (modulo an additive constant). In other words, such GFF preserves shift symmetry. Hence, we can compactify the target space and define a compact version of such GFF on the space of continuous functions $\mathbb{R}\rightarrow S^{1}_{1/b_{0}}$.
@@
Due to the identification in field space, all well-defined observables must be built out of $\partial^{n}_{x}\phi(x)$ with $n\geq 1$, and vertex operators $e^{\mathrm{i}nb_{0}\phi(x)}$, with $n\in\mathbb{Z}$. The correlators of $\partial^{n}_{x}\phi(x)$ are obtained straightforwardly as derivatives of (products of) $C(x)$, and so they are clearly independent of the additive constant; those of the vertex operators will be discussed below. All such operators have positive dimension, and the resulting theory is unitary, as proven rigorously in [58].131313 More precisely, in [58] the authors proved reflection positivity of the fractional Brownian motion, that is obtained from the GFF $\phi$ by fixing its value to be zero at the origin.
We point out that while this GFF might seem a bit exotic, its status is not much different from the standard massless scalar field in $d=1$. The latter corresponds to taking $\delta=1$, and thus has negative scaling dimension. Therefore, it also needs to be defined by modding out its zero mode, and thus also in this case the well-defined random variables are those that are independent of the zero mode. In particular, the variable $\phi(x)-\phi(0)$ corresponds to the standard Wiener process (Brownian motion), and the growth of the covariance reproduces the well-known linear growth of the Brownian motion: $\langle(\phi(x)-\phi(0))^{2}\rangle_{0}\sim|x|$. The generalization to $\delta\neq 1$ is known as fractional Brownian motion, and it was introduced long ago in [59]. Lastly, the compactification simply corresponds to the restriction of the random process to a circle.
@@
Since $C(x)$ is defined only up to a constant, we could set $C(0)$ to zero. However, for the $\delta\to 0$ limit we choose $C(x)$ as in (3.2), and thus absorb $C(0)$ by a normal ordering of the vertex operators, consistently with the $\delta\to 0$ limit, where such normal ordering is needed because of UV divergences (see footnote 14). We thus define the vertex operators at $\delta>0$ as:
@@ 3.2 Interacting model
The first justification of our claim is that we reproduce the AYK partition function (2.20) by expanding (3.9) in $g$. In order to see that, let us first set $h=0$, and expand $Z_{s}(b_{0},g)$ in powers of $g$. Using (3.12), with $A=\mathcal{O}_{g}$, we find precisely (2.20), upon the identification $b_{0}^{2}=\mathcal{J}$. Indeed, since $\hat{\sigma}_{+}^{2}=\hat{\sigma}_{-}^{2}=0$, only alternating chains of kink and antikink operators $\hat{\sigma}_{+}V_{+}(x)$ and $\hat{\sigma}_{-}V_{-}(x)$ survive. Turning on $h$, we will show below that it is equivalent to shifting $b_{0}$ to $b_{0}-\sqrt{2}h$, hence the partition function only depends on the combination $b=b_{0}-\sqrt{2}h$ and the correspondence with AYK still holds, with $\mathcal{J}=b^{2}$. However, keeping the $\mathcal{O}_{h}$ term will prove essential in order to consistently renormalize the model.
The partition function (3.9) is similar to those encountered in impurity models [63, 64], and the path ordering of $2\times 2$ matrices can be traded for a path integral using a complex bosonic spinor $z(x)=\{z_{1}(x),z_{2}(x)\}$ subject to $\bar{z}(x)z(x)=1$, similarly to [65, 63, 64]. We will review this formulation in Appendix C.1. See also Appendix C.3 for an alternative formulation in terms of a nonlinear sigma model.
@@
For $s=1-\delta<1$, the model (3.9) is a genuinely new theory. It might seem rather unorthodox, since $\Delta_{\phi}<0$ and since the vertex operators do not have a definite scaling dimension in the UV. Nevertheless, it is consistent, as it should be be clear from the previous discussion of the GFF with negative dimension, and from the analysis presented in the rest of the paper. [^16] We notice moreover that similar type of non-polynomial non-scaling operators appears also in other contexts, such as in Lagrangians of nonlinear sigma models and in low-energy effective actions of QED3 [67, 68].
@@ 3.2.1 Relation between $h$ and $b$
Therefore, up to a boundary contribution, [^17] If we had defined the model on a circle, i.e. with periodic boundary conditions at x=L/2x=-L/2 and $x=L/2$, the factor $\mathcal{D}_{h}(L/2,-L/2)$ would be trivial if $\phi$ was non-compact, while it would give a nontrivial monodromy in the compact case. we find that the effect of the new interaction term is equivalent to shifting $b_{0}\to b\equiv b_{0}-h\sqrt{2}$, and renormalizing the coupling $g$. This implies that we can trade a change in $b_{0}$ for a change in $h$, or viceversa. This fact is often used in the Kondo model (i.e. at $\delta=0$) in order to set h=0h=0.181818The fact that one can trade $b$ for $h$, or vice versa, is mentioned in various old papers about the Kondo model, but typically it is either stated without explanation (e.g. [22, 53]) or expressed as a canonical transformation on the quantum Hamiltonian (e.g. [66, 69, 70]). We have not found a proof in the literature that uses the path integral language as does the one provided here. We will instead set $b_{0}$ to our favorite value (in particular, such that $b_{0}=1$ when $\delta=0$) and keep $h\neq 0$ in our model. As we will see in Section 4, the choice of keeping a nonvaninshing $h$ is forced upon us by the renormalization flow: indeed the choice $h=0$ is not stable under renormalization.

### v2

@@ 1 Introduction and motivation
While a simple-minded generalization of (1.3) – coupling the generalized free field $\chi$ (with $\Delta_{\chi}=(1+s)/2$) to the topological operator $\sigma$ (with $\Delta_{\sigma}=0$) via $\sigma\chi$ – would correctly identify the crossover location at $s^{\star}=1$, it would fail to provide a correct description of the critical theory. Indeed, $\sigma$ could be represented in the standard qubit basis as the third Pauli matrix $\hat{\sigma}_{3}$, and it is easily checked that it would never become a genuine local field, it would not acquire an anomalous dimension, and the relevant operator controlling the deviation from criticality would also absent. [^1] These remarks might become more transparent after reading section 3, and noticing that the naive model discussed here would correspond to the $g=0$ case of our complete model. A different perspective is therefore needed in order to capture the full operator spectrum and scaling behavior near the crossover to short range in one dimension.
@@
  • For all $s\leq 1$, we propose a field theory that is weakly coupled near the crossover, exactly solvable at $s=1$, that reduces to the Kondo model at $s=1$, and that upon perturbative expansion reproduces the AYK model. At $s>1$, it reproduces the 1d SRI physics, via a trivial fixed point, reached only at zero temperature. We derive several predictions for CFT data at next-to-leading or next-to-next-to-leading order in perturbation theory, thus showing that such model provides not only a conceptual framework to recast the AYK model as a field theory, but also a computational tool to treat the near-crossover regime with systematic perturbative methods.
@@ 2.2.1 Unitary 1d CFTs with parity symmetry: the bootstrap definition
The lowest-weight state in $D_{\Delta_{i},J_{i}}$ corresponds to a local primary operator $\phi_{i}$ at $x=0$. $\Delta_{i}$ is the scaling dimension of $\phi_{i}$, and $J_{i}\in\{0,1\}$ its parity, so that $x\mapsto-x$ acts as $\phi_{i}(0)\mapsto(-1)^{J_{i}}\phi_{i}(0).$ The sequence of ordered pairs $((\Delta_{i},J_{i}))_{i=0}^{\infty}$ is called the spectrum of the theory. Reflection positivity implies $\Delta_{i}\geq 0$. We have $\Delta_{0}=0$, $J_{0}=0$, corresponding to the vacuum (in other words, $\phi_{0}$ is identity operator $\mathds{1}$). Correlation functions of $\phi_{i}$ are invariant under $\mathrm{PGL}_{2}(\mathbb{R})$, which maps point $x$ to $\frac{ax+b}{cx+d}$ and the field $\phi$ to $\phi^{\prime}$, given by [^6] Note that $\tfrac{dx-b}{-cx+a}$ is the inverse of $\tfrac{ax+b}{cx+d}$. Indeed, when a symmetry $g$ acts on a space by $x\mapsto gx$, the induced action on functions sends $f$ to $f^{\prime}(x)=f(g^{-1}x)$. |cx+a||-cx+a| is defined using a representative $g\in\mathrm{GL}_{2}(\mathbb{R})$ of $g]\in\mathrm{PGL}_{2}(\mathbb{R})$ with $|\mathrm{det}(g)|=1$.
@@
holding for all $i,j,k,\ell\in\mathbb{Z}_{>0}$ and all $0<z<1$. Here $G^{\Delta_{i},\Delta_{j},\Delta_{k},\Delta_{\ell}}_{\Delta_{m}}(z)$ are the 1d conformal blocks, computed in [40, 41, 42], taking the form
@@ 2.3 Anderson-Yuval-Kosterlitz model
Figure 3: The RG flow for the beta functions in (2.21), at $s=1$ and $s=0.9$. Arrows point towards the IR. At $s=1$, the $g=0$ axis is a line of fixes points, but we have emphasized in blue and red the special points $\mathcal{J}=0$ and $\mathcal{J}=1$, respectively. The red lines separate the diagram into three or four regions: region II is the low-temperature broken phase ($g=0$ in the IR); region IIII is the high-temperature symmetric phase ($g\sim O(1)$ in the IR); regions $III$ and $IV$ are inaccessible to the LRI. The latter indeed corresponds to a one-dimensional subspace such as the dashed blue line (see (B.6)). The intersection of such line with the red line defines the critical coupling $\mathcal{J}_{c}$ (i.e. critical temperature at fixed bond strength $J$, if $\mathcal{J}=J/T$).
@@ 2.4 Insights from the 1d short-range Ising model
Another important insight for developing a theory of the LRI model near $s=1$ comes from scrutinizing the possible physical content of the theory at $s>1$. As reviewed above, for $d\geq 2$ and $s\geq s^{*}$, the LRI crosses to the SRI universality plus a decoupled GFF sector [10, 9]. What changes when $d=1$? As we argue below, the physics is bound to be not so different from that of the classical 1d SRI model, indeed:
@@
Given the facts we just recalled about the SRI, we would expect that at $s\geq 1$ the critical LRI, with classical Hamiltonian (1.1), would be described by the degrees of freedom of a $\mathbb{C}^{2}$ Hilbert space, with trivial correlators. However, as in $d>1$, this cannot be the end of the story, because it would imply that at the crossover some operators would disappear from the spectrum. For example, by continuity of spectrum, at $s=1$ we should find two $\mathbb{Z}_{2}$-odd operators with opposite parity and of dimension exactly equal to one, corresponding to $\varphi^{3}$ and $\partial\varphi$ in the $\varphi^{4}$ description (see Section 2.2), but clearly the SRI alone cannot provide such operators. [^11] Following [9, 10], we could try to account for the missing operators by adding to the 1d SRI a GFF $\chi$ of scaling dimension $\Delta_{\chi}=(1+s)/2$, that could be identified with $\varphi^{3}$, and that should be decoupled from the spin degree of freedom at $s\geq 1$. However, we would still be missing the parity-odd operator corresponding to $\partial\varphi$, as it would be unclear how to introduce derivatives of $\hat{\sigma}_{3}$. Moreover, it would be unclear how to couple the two sectors at $s<1$ in order to flow to the LRI CFT: if for example, combining insights from the above reminder of 1d SRI and from [9, 10], we introduce the defect operator $\operatorname{tr}[\exp(h\,\int{\rm d}x\,\hat{\sigma}_{3}\,\chi(x))]$ in correlators over the $\mathbb{C}^{2}+$GFF theory, we obtain a trivial result, as the trace vanishes.
@@ 3.1 Compact generalized free fields with negative scaling dimension
As we review in Appendix A, for $1<\Delta_{\phi}\leq 0$, the noncompact 1d GFF can be rigorously defined [58] as a probability measure on the space of distributions $\mathbb{R}\rightarrow\mathbb{R}$ defined modulo an additive constant. In fact, the resulting random distribution is almost surely a continuous function modulo an additive constant. In other words, such GFF preserves shift symmetry. Hence, we can compactify the target space and define a compact version of such GFF on the space of continuous functions $\mathbb{R}\rightarrow S^{1}_{1/b_{0}}$.
@@
Due to the identification in field space, all well-defined observables must be built out of $\partial^{n}_{x}\phi(x)$ with $n\geq 1$, and vertex operators $e^{\mathrm{i}nb_{0}\phi(x)}$, with $n\in\mathbb{Z}$. The correlators of $\partial^{n}_{x}\phi(x)$ are obtained straightforwardly as derivatives of (products of) $C(x)$, and so they are clearly independent of the additive constant; those of the vertex operators will be discussed below. All such operators have positive dimension, and the resulting theory is unitary, as proven rigorously in [59]. [^14] More precisely, in [59] the authors proved reflection positivity of the fractional Brownian motion, that is obtained from the GFF $\phi$ by fixing its value to be zero at the origin.
Remark 1.
We point out that while this GFF might seem a bit exotic, its status is not much different from the standard massless scalar field in $d=1$. The latter corresponds to taking $\delta=1$, and thus has negative scaling dimension. Therefore, it also needs to be defined by modding out its zero mode, and thus also in this case the well-defined random variables are those that are independent of the zero mode. In particular, the variable $\phi(x)-\phi(0)$ corresponds to the standard Wiener process (Brownian motion), and the growth of the covariance reproduces the well-known linear growth of the Brownian motion: $\langle(\phi(x)-\phi(0))^{2}\rangle_{0}\sim|x|$. The generalization to $\delta\neq 1$ is known as fractional Brownian motion, and it was introduced long ago in [60]. Lastly, the compactification simply corresponds to the restriction of the random process to a circle, and in the $\delta=1$ case it is known as $O(2)$ rigid rotator in the physics literature (e.g. [61]).
Remark 2.
Let us ignore for a moment the vertex operators, and define $\chi(x)\sim\partial_{x}\phi(x)$. With the measure defined by the expectation value $\langle\cdot\rangle_{0}$, such field $\chi$ together with its derivatives and their composite (normal ordered) operators form a set of operators of a CFT (as defined in Section 2.2.1), closed under OPE. Clearly such CFT is nothing else than the GFF $\chi$ of conformal dimension $\Delta_{\chi}=1-\delta/2$. In this case, the compactification of $\phi$ has no role and it would be more logical to directly start with the GFF $\chi$. However, as will become clear below, in order to construct a dual theory of the 1d LRI we need the vertex operators. These would be nonlocal operators when written in terms of $\chi$, and although similar nonlocal operators appear for example when writing correlators of the XY model in terms of a sine-Gordon field [62], it seems more natural to work with the ancestor of $\chi$, namely the field $\phi$.
Remark 3.
Once we include the vertex operators in our set of operators, the GFF $\phi$ is no longer a CFT, because at $\delta>0$ their correlators are not conformal and they depend explicitly on the dimensionful compactification radius $1/b_{0}$ (see below). Although the theory is Gaussian, the presence of a dimensionful parameter leads to a classical renormalization group flow, that is, introducing a renormalization scale $\mu$ and defining $\tilde{b}_{0}=\mu^{-\delta/2}b_{0}$, we have a linear beta function $\beta_{\tilde{b}_{0}}=-\frac{\delta}{2}\tilde{b}_{0}$. This has a UV fixed point at $\tilde{b}_{0}=0$, where the vertex operators become degenerate with the identity, and we recover the CFT of the previous remark. Loosely speaking, we might qualitatively compare the situation to that of a standard free massive scalar (with the mass playing the role of $b_{0}$), which likewise becomes a CFT only in the UV limit.
@@
Since $C(x)$ is defined only up to a constant, we could set $C(0)$ to zero. However, for the $\delta\to 0$ limit we choose $C(x)$ as in (3.2), and thus absorb $C(0)$ by a normal ordering of the vertex operators, consistently with the $\delta\to 0$ limit, where such normal ordering is needed because of UV divergences (see footnote 15). We thus define the vertex operators at $\delta>0$ as:
@@ 3.2 Interacting model
The first justification of our claim is that we reproduce the AYK partition function (2.20) by expanding (3.9) in $g$. In order to see that, let us first set $h=0$, and expand $Z_{s}(b_{0},g)$ in powers of $g$. Using (3.12), with $A=\mathcal{O}_{g}$, we find precisely (2.20), upon the identification $b_{0}^{2}=\mathcal{J}$. Indeed, since $\hat{\sigma}_{+}^{2}=\hat{\sigma}_{-}^{2}=0$, only alternating chains of kink and antikink operators $\hat{\sigma}_{+}V_{+}(x)$ and $\hat{\sigma}_{-}V_{-}(x)$ survive. Turning on $h$, we will show below that it is equivalent to shifting $b_{0}$ to $b_{0}-\sqrt{2}h$, hence the partition function only depends on the combination $b=b_{0}-\sqrt{2}h$ and the correspondence with AYK still holds, with $\mathcal{J}=b^{2}$. However, keeping the $\mathcal{O}_{h}$ term will prove essential in order to consistently renormalize the model, as we will show in section 4. The fact that such field theoretic renormalization based on (3.9) will allow us to reproduce (and go beyond) the beta functions (2.21) will provide a second justification of our claim. Further support, such as the matching of symmetries to those of the LRI and the fulfillment of constraints on the spectrum of protected operators and on OPE coefficients, which are characteristic of the LRI, will be detailed in the following.
The partition function (3.9) is similar to those encountered in impurity models [66, 67], and the path ordering of $2\times 2$ matrices can be traded for a path integral using a complex bosonic spinor $z(x)=\{z_{1}(x),z_{2}(x)\}$ subject to $\bar{z}(x)z(x)=1$, similarly to [68, 66, 67]. We will review this formulation in Appendix C.1. One advantage of such “coherent state representation” is that the Pauli matrices are replaced by fields, thus giving the model a more standard field theoretic appearance, and making it easier to derive Schwinger-Dyson equations and introduce a gauging (as we do in Appendix C.2). On the other hand, the “Pauli representation” is more convenient for doing perturbative computations, and it makes the link to the SRI more transparent.
See also Appendix C.3 for an alternative formulation in terms of a (long-range) nonlinear sigma model. The advantage of this other reformulation is to provide a different point of view on the compactness of $\phi$, and to allow a more standard implementation of the gauging in the GFF sector. However, for practical computations we are forced to solve the nonlinear constraints and thus reintroduce $\phi$.
@@
For $s=1-\delta<1$, the model (3.9) is a genuinely new theory. It might seem rather unorthodox, since $\Delta_{\phi}<0$ and since the vertex operators do not have a definite scaling dimension in the GFF. However, it should be clear from the discussion above that the negative-dimension GFF is a well-defined QFT, and we will see from the RG analysis (in particular end of section 4.3.1) that qualitatively the situation with the perturbation is not too different from that of a standard massive free scalar perturbed by self-interactions, which are also not scaling operators for a massive theory. [^17] We notice moreover that similar types of non-polynomial non-scaling operators appear also in other contexts, such as in Lagrangians of nonlinear sigma models and in low-energy effective actions of QED3 [70, 71]. Moreover, the analysis presented in the rest of the paper should provide further evidence that it is a fully consistent theory.
@@ 3.2.1 Relation between $h$ and $b$
Therefore, up to a boundary contribution, [^18] If we had defined the model on a circle, i.e. with periodic boundary conditions at x=L/2x=-L/2 and $x=L/2$, the factor $\mathcal{D}_{h}(L/2,-L/2)$ would be trivial if $\phi$ was non-compact, while it would give a nontrivial monodromy in the compact case. we find that the effect of the new interaction term is equivalent to shifting $b_{0}\to b\equiv b_{0}-h\sqrt{2}$, and renormalizing the coupling $g$. This implies that we can trade a change in $b_{0}$ for a change in $h$, or viceversa. This fact is often used in the Kondo model (i.e. at $\delta=0$) in order to set h=0h=0.191919The fact that one can trade $b$ for $h$, or vice versa, is mentioned in various old papers about the Kondo model, but typically it is either stated without explanation (e.g. [22, 54]) or expressed as a canonical transformation on the quantum Hamiltonian (e.g. [69, 72, 73]). We have not found a proof in the literature that uses the path integral language as does the one provided here. We will instead set $b_{0}$ to our favorite value (in particular, such that $b_{0}=1$ when $\delta=0$) and keep $h\neq 0$ in our model. As we will see in Section 4, the choice of keeping a nonvaninshing $h$ is forced upon us by the renormalization flow: indeed the choice $h=0$ is not stable under renormalization.